# Exposing Lip-syncing Deepfakes from Mouth Inconsistencies

### v1

@@ 4.2 Results
Cross-type deepfakes testing. Furthermore, we assess the performance of our model on three face-swap datasets, using the model trained on the FakeAV-LS set. The LIPINC model achieves an AUC score of 91% on the FakeAV-FS dataset as in-domain cross-type testing, while an AUC of 60% and 67% on the KODF-FSGAN and KODF-DFL [24] datasets, respectively. The drop in performance on the KODF dataset suggests the presence of distinct mouth inconsistency patterns between lip-syncing videos and face-swap deepfakes.
@@
Ablation studies. In Fig.4, we show the ablation studies to validate the efficacy of different modules in our model on the two datasets, namely LSR+W2L and KODF-LS [24]. Our proposed LIPINC utilizes five local frames, three global frames, color and structure branches in the MSTIE module, and an inconsistency loss to detect deepfakes. Here C_w/o_S+MSTIE refers to the model with only the color branch without the MSTIE module. Similarly, S_w/o_C+MSTIE refers to the model with only a structure branch. C+S_w/o_MSTIE indicates a model with two branches without the MSTIE module, and C+S_w/o_IL indicates a model without our inconsistency loss. We use L for the local frames and G for the global frames. The associated number refers to the number of frames for that case. The comparison results in Fig.4 clearly demonstrate the contribution of each module in our model architecture and how variations in a number of local or global frames influence the performance of our model.
@@
Specifically, it can be observed that the removal of any module results in an obvious decrease in the generalization ability of the model. First, removing all the local frames (3_G) affects the performance the most, where the AUC decreases by 14% and 25% on the LSR+W2L and KODF_LS, respectively, whereas the removal of the global frames (5_L) reduced the AUC by 4% and 2% on two datasets respectively. Second, removing the inconsistency loss (C+S_w/o_IL) decreased the AUC by 17% for both datasets. Similarly, when we delete the MSTIE module, the performance of C_w/o_S+MSTIE decreases by 26% and 15%, respectively. However, it should be noted that the performance of S_w/o_C+MSTIE remains adequate, with the AUC decreasing by just 2% for both LSR+W2L and KODF-LS, showing the importance of structure features in inconsistency learning. Interestingly, when the color branch is added back to the S_w/o_C+MSTIE model, namely the C+S_w/o_MSTIE strategy, the performance deteriorates. This happens because using the concatenate layer instead of our MSTIE to fuse the outputs of the color and structure branch cannot effectively capture the inconsistencies. Last, we find that decreasing the number of the total frames slightly affects the generalization performance. It can be noted that the reduction of global frames from 3 to 2 (5_L+2_G) affected the performance more than the reduction of local frames from 5 to 3 (3_L+3_G).
@@
Limtations. Although our model performs well in detecting lip-syncing deepfakes, our approach cannot detect deepfake images and videos with no apparent lips, or videos that are too short to detect global frames.
5 Conclusion
@@
In this paper, we have proposed a novel method called LIPINC to detect lip-syncing deepfakes. We demonstrate that the mouth region in deepfake videos contains spatial-temporal discrepancies. These inconsistencies are obvious in the consecutive frames and throughout the video. Our model can capture these inconsistency patterns by designing a novel mouth spatial-temporal inconsistency extractor guided by an inconsistency loss. Extensive experimental results on three datasets show the outstanding performance of our method in detecting both in-domain and cross-domain deepfakes.
Acknowledgements: This work is supported by the US Defense Advanced Research Projects Agency (DARPA) Semantic Forensic (SemaFor) program, under Contract No. HR001120C0123.

### v2

@@ IV-D Cross-type Deepfake Testing
Furthermore, to demonstrate the performance of our LIPINC model across different types of deepfake videos, such as faceswap deepfakes, we evaluate its performance on three face-swap datasets using the model trained on the FakeAV-LS set. The LIPINC model achieves an AUC score of 91% on the FakeAV-FS dataset as in-domain cross-type testing, while an AUC of 60% and 67% on the KODF-FSGAN and KODF-DFL [23] datasets, respectively. The drop in performance on the KODF dataset suggests the presence of distinct mouth inconsistency patterns between lip-syncing videos and face-swap deepfakes.
IV-E Ablation Studies
@@
In this section, we conduct ablation studies to validate the efficacy of different modules in our model on the two datasets, namely LSR+W2L and KODF-LS [23]. Our proposed LIPINC model utilizes five local frames, three global frames, color and structure branches in the MSTIE module, and an inconsistency loss to detect deepfakes. We consider several variants of our detection model, including C_Only which refers to the model with only the color branch without the MSTIE module, S_Only which refers to the model with only a structure branch, C+S_Concat indicating a model with two branches without the MSTIE module, and C+S_w/o_IL indicating a model without our inconsistency loss. We use L for the local frames and G for the global frames. The associated number refers to the number of frames for that case. The comparison results are shown in Fig. 4. It clearly demonstrates the contribution of each module in our model architecture and how variations in a number of local or global frames influence the performance of our model.
@@
Specifically, it can be observed that the removal of any module results in an obvious decrease in the generalization ability of the model. First, removing all the local frames (3_G) affects the performance the most, where the AUC decreases by 14% and 25% on the LSR+W2L and KODF_LS, respectively, whereas the removal of the global frames (5_L) reduced the AUC by 4% and 2% on two datasets respectively. Second, removing the inconsistency loss (C+S_w/o_IL) decreased the AUC by 17% for both datasets. Similarly, when we delete the MSTIE module, the performance of C_Only decreases by 26% and 15%, respectively. However, it should be noted that the performance of S_Only remains adequate, with the AUC decreasing by just 2% for both LSR+W2L and KODF-LS, showing the importance of structure features in inconsistency learning. Interestingly, when the color branch is added back to the S_Only model, namely the C+S_Concat strategy, the performance deteriorates. This happens because using the concatenate layer instead of our MSTIE to fuse the outputs of the color and structure branch cannot effectively capture the inconsistencies. Last, we find that decreasing the number of the total frames slightly affects the generalization performance. It can be noted that the reduction of global frames from 3 to 2 (5_L+2_G) affected the performance more than the reduction of local frames from 5 to 3 (3_L+3_G).
IV-F Limitations
@@
Although our model performs well in detecting lip-syncing deepfakes, our approach cannot detect deepfake images and videos lacking lip movement, or videos that are too short to detect global frames.
V Conclusion
@@
In this paper, we have proposed a novel method called LIPINC to detect lip-syncing deepfakes. We demonstrate that the mouth region in deepfake videos contains spatial-temporal discrepancies. These inconsistencies are obvious in the consecutive frames and throughout the video. Our model can capture these inconsistency patterns by designing a novel mouth spatial-temporal inconsistency extractor guided by an inconsistency loss. Extensive experimental results on three datasets show the outstanding performance of our method in detecting both in-domain and cross-domain deepfakes. Our future work will explore incorporating the audio modality to enhance lip-syncing detection performance.
Acknowledgement. This work is supported in part by the US Defense Advanced Research Projects Agency (DARPA) Semantic Forensic (SemaFor) program, under Contract No. HR001120C0123, National Science Foundation (NSF) Projects under grants SaTC-2153112, No.1822190, and TIP-2137871, and University at Buffalo’s Office of Vice President for Research and Economic Development and Center for Information Integrity. The views and conclusions contained herein are those of the authors and should not be interpreted as necessarily representing the official policies, either expressed or implied, of DARPA, NSF, or the U.S. Government.